# Directional Sheaf Hypergraph Networks: Unifying Learning on Directed and Undirected Hypergraphs

### v1

@@ 5 Experimental evaluation
We evaluate DSHN and DSHNLight against 13 baseline models from the directed and undirected hypergraph literature on real-world datasets (Section 5.1) as well as synthetic ones (Section 5.2) on the node classification task. Model performance is measured in terms of classification accuracy. Following the standard practice in the literature (Chien et al., 2022; Wang et al., 2023a; Fiorini et al., 2024), we adopt a 50%/25%/25% split for training, validation, and testing, respectively, and, for each model, we report the average test accuracy and the standard deviation over 10 independent runs. Details on the baselines, hyperparameter tuning, and the experimental setup are provided in Appendix C.
@@ 5.1 Real-World Datasets
To evaluate our models on real-world datasets, we follow the pre-processing procedure introduced by Tran and Tran (2022) and Fiorini et al. (2024), and apply it to a suite of publicly available directed graph benchmarks to obtain their directed hypergraph counterparts for performing the node classification task (see Section C.5). The considered datasets are: Cora (Zhang et al., 2022), email-Enron, email-EU (Benson et al., 2018b), Telegram (Bovet and Grindrod, 2020), Chameleon, Squirrel, and Roman-empire. Due to space limitations, Table 1 includes only the datasets that yield the most interesting insights. Additional results can be found in Table 3, while additional informations on the datasets are provided in Section C.4.
@@
DSHN, and its variant DSHNLight, which both leverage the theoretical advantages of associating a Directed Cellular Sheaf to a directed hypergraph, consistently outperform the 13 baselines from both the undirected and directed hypergraph learning literature on 6 out of 7 real-world datasets. The largest relative gains are observed on the email-Enron and email-EU datasets, where DSHN and DSHNLight improve over the best baseline by more than 20%. A substantial improvement is also achieved on Telegram, confirming the importance of directional information in this benchmark where all directed methods perform strongly. More moderate but consistent improvements are found on highly heterophilic datasets such as Roman-empire, Chameleon, and Squirrel, while on highly homophilic datasets such as Cora performance is on par with the strongest baselines. As shown in Table 2, the charge parameter $q$ selected by the hyperparameter-selection procedure for our models on highly homophilic datasets is consistently $0.0$. This observation is in line with the findings of Zhang et al. (2021), who report that, in such settings, directional information behaves as noise for node classification.
@@ 5.2 Synthetic Datasets
We additionally evaluate our models on the synthetic datasets introduced by Fiorini et al. (2024), built over $n=500$ nodes and split into $c=5$ classes. Each class contains 30 random intra-class hyperedges, while inter-class directed hyperedges, consisting of multiple tail and head nodes, are added between class pairs with sizes drawn uniformly from $\{3,\ldots,10\}$. By varying the number of inter-class hyperedges $I_{o}\in\{10,30,50\}$, we control the strength of directional connectivity. This design provides a clean benchmark to test the models’ ability to capture directionality; further details are given in Section C.4.
@@
The results in Table 3 clearly demonstrate the advantage of our models DSHN and DSHNLight over existing baselines. Classical undirected hypergraph methods are unable to capture the directional structure that dominate these benchmarks, and as a result their performance is limited. Directed methods such as GeDi-HNN and DHGNN achieve stronger results, confirming the importance of explicitly incorporating directionality into the convolutional process. Yet, DSHN and DSHNLight, which provide a principled and more expressive treatment of directional structure, yield consistent improvements across all synthetic datasets, outperforming the best directed baselines by up to 18 percentage points and reaching 99.04% accuracy on the third synthetic dataset—this highlights the expressive power that the notion of Directed Hypergraph Cellular Sheaves unlocks.
@@ 6 Conclusion and Future Works
In this work, we introduced the concept of Directed Hypergraph Cellular Sheaves for directed hypergraphs and derived the corresponding Directed Sheaf Hypergraph Laplacian, which we integrated into our proposed framework DSHN. By encoding hyperedge direction via a topology-aware complex-valued inductive bias, our method naturally accommodates both directed and undirected hypergraphs while also unifying and generalizing several operators from the graph and hypergraph learning literature.
Across a broad set of benchmark datasets, DSHN consistently outperforms methods from both the directed and undirected hypergraph learning literature. As future work, a natural step forward is to evaluate our framework on larger and natively directed hypergraph datasets such as protein-protein interaction networks to further test the scalability and expressivity of the method. Finally, an intriguing direction is to make the charge parameter $q$ directly learnable, allowing each layer to adapt its diffusion process dynamically.
@@ A.1 Spectral Properties
The following Lemma (which we state for clarity even if it is not reported in the paper as a lemma) derives the expression of our proposed Laplacian matrix $L^{\vec{\mathcal{F}}}$ when applied as a linear operator on a signal:
@@
Let $x\in\mathbb{C}^{nd}$ be a complex-valued signal. Component-wise, the application of $L^{\vec{\mathcal{F}}}$ to it and to its normalized counterpart reads:
$\bigl(L^{\vec{\mathcal{F}}}(x)\bigr)_{u}=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,x_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,x_{v}\right).$
$\bigl(L^{\vec{\mathcal{F}}}_{N}(x)\bigr)_{u}=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\Bigl(D_{u}^{-\frac{1}{2}}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\Bigr)\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,D_{u}^{-\frac{1}{2}}x_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,D_{v}^{-\frac{1}{2}}x_{v}\right).$
@@
We start by applying the definition of the $L^{\vec{\mathcal{F}}}$ component-wise as in Eq. 3:
$\bigl(L^{\vec{\mathcal{F}}}(x)\bigr)_{u}$ $=$ $\sum_{v\in V}(L^{\vec{\mathcal{F}}})_{uv}\,x_{v}$
$=$ $\sum_{e:\,u\in e}\left(1-\frac{1}{\delta_{e}}\right)\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{u\trianglelefteq e}\,x_{u}-\sum_{e:\,u\in e}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\,x_{v}$
$=$ $\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\left((\delta_{e}-1)\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{u\trianglelefteq e}\,x_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\,x_{v}\right)$
$=$ $\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\left((\delta_{e}-1)\,\vec{\mathcal{F}}_{u\trianglelefteq e}\,x_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\vec{\mathcal{F}}_{v\trianglelefteq e}\,x_{v}\right).$
Finally, notice that the coefficient $\delta_{e}-1$ is exactly the number of vertices in $e$ different from $u$. Thus, the term $(\delta_{e}-1)\,\vec{\mathcal{F}}_{u\trianglelefteq e}\,x_{u}$ can be written as a sum of $\vec{\mathcal{F}}_{u\trianglelefteq e}\,x_{u}$ over all $v\in e,v\neq u$. Substituting this back, we obtain:
$\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\left(\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\vec{\mathcal{F}}_{u\trianglelefteq e}\,x_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\vec{\mathcal{F}}_{v\trianglelefteq e}\,x_{v}\right)$
$\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,x_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,x_{v}\right).$
@@
The claim follows rather directly since, as it is not hard to see, $L^{\vec{\mathcal{F}}}_{N}$ is Hermitian by construction. ∎
@@
By definition of the Dirichlet energy as the quadratic form associated with $L^{\vec{\mathcal{F}}}_{N}$, we have:
$\mathcal{E}_{N}(x)=x^{\dagger}L^{\vec{\mathcal{F}}}_{N}x=\sum_{u\in V}x_{u}^{\dagger}\bigl(L^{\vec{\mathcal{F}}}_{N}(x)\bigr)_{u}.$
By substituting for $(L^{\vec{\mathcal{F}}}_{N}(x))_{u}$ (see the previous lemma), we have:
$\mathcal{E}_{N}(x)$ $=\sum_{u\in V}\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr)^{\dagger}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,D_{u}^{-\tfrac{1}{2}}x_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,D_{v}^{-\tfrac{1}{2}}x_{v}\right).$
@@
$\mathcal{E}_{N}(x)$ $=\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{u\in e}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}-\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}$
$=\sum_{e\in E}\frac{1}{\delta_{e}}\left(\sum_{u\in e}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\bigl\|\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr\|_{2}^{2}-\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}\right).$
Since $L^{\vec{\mathcal{F}}}_{N}$ is Hermitian, the second inner summation can be rewritten as:
$-\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}=$
$-\sum_{\begin{subarray}{c}u,v\in e\\ u<v\end{subarray}}\left((\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}+\bigl(\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}\bigr)^{\dagger}\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\right)=$
$-\sum_{\begin{subarray}{c}u,v\in e\\ u<v\end{subarray}}2\,\Re\left[\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}\right]=$
$-\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\Re\left[\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}\right].$
@@
$\mathcal{E}_{N}(x)$ $=\frac{1}{2}\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\Bigl(\|\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\|_{2}^{2}+\|\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}\|_{2}^{2}-2\,\Re\left[\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}\right]\Bigr).$
@@
$\mathcal{E}_{N}(x)=\frac{1}{2}\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\Bigl\|\vec{\mathcal{F}}_{u\trianglelefteq e}D_{u}^{-\tfrac{1}{2}}x_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}D_{v}^{-\tfrac{1}{2}}x_{v}\Bigr\|_{2}^{2}.$

### v2

@@ 5 Experimental Evaluation
We evaluate DSHN and DSHNLight against 13 baseline models from both the directed and undirected hypergraph literature on real-world datasets (Section 5.1) as well as synthetic datasets (Section 5.2) for the node classification task. From the undirected hypergraph-learning literature, we include HGNN (Feng et al., 2019), HNHN (Dong et al., 2020), UniGCNII (Huang and Yang, 2021), LEGCN (Yang et al., 2022), HyperND (Tudisco et al., 2021), AllDeepSets and AllSetTransformer (Chien et al., 2022), ED-HNN (Wang et al., 2023a), SheafHyperGNN (Duta et al., 2023), and PhenomNN (Wang et al., 2023b). From the directed hypergraph-learning literature, we consider GeDi-HNN (Fiorini et al., 2024) and DHGNN (Ma et al., 2024), along with a variant, as baselines. Model performance is measured in terms of classification accuracy. Following the standard practice in the literature (Chien et al., 2022; Wang et al., 2023a; Fiorini et al., 2024), we adopt a 50%/25%/25% split for training, validation, and testing, respectively, and, for each model, we report the average test accuracy and the standard deviation over 10 independent runs. Details on the baselines, hyperparameter tuning, and the experimental setup are provided in Appendix D.
@@ 5.1 Real-World Datasets
To evaluate our models on real-world datasets, we follow the pre-processing procedure introduced by Tran and Tran (2022) and Fiorini et al. (2024), and apply it to a suite of publicly available directed graph benchmarks to obtain their directed hypergraph counterparts for performing the node classification task (see Section D.5). The considered datasets are: Cora (Zhang et al., 2022), email-Enron, email-EU (Benson et al., 2018b), Telegram (Bovet and Grindrod, 2020), Chameleon, Squirrel, and Roman-empire. Due to space limitations, Table 1 includes only the datasets that yield the most interesting insights. Additional results can be found in Table 3, while additional informations on the datasets are provided in Section D.4.
@@
DSHN, and its variant DSHNLight, which both leverage the theoretical advantages of associating a Directed Cellular Sheaf to a directed hypergraph, consistently outperform the 13 baselines from both the undirected and directed hypergraph learning literature on 6 out of 7 real-world datasets. The largest relative gains are observed on the email-Enron and email-EU datasets, where DSHN and DSHNLight improve over the best baseline by up to 20%. A substantial improvement is also achieved on Telegram, confirming the importance of directional information in this benchmark where all directed methods perform strongly. More moderate but consistent improvements are found on highly heterophilic datasets such as Roman-empire, Chameleon and Squirrel while on highly homophilic datasets such as Cora performance is on par with the strongest baselines. As shown in Table 2, the charge parameter $q$ selected by the hyperparameter-selection procedure for our models on highly homophilic datasets is consistently $0.0$. This observation is in line with the findings of Zhang et al. (2021), who report that, in such settings, directional information behaves as noise for node classification.
@@ 5.2 Synthetic Datasets
We additionally evaluate our models on the synthetic datasets introduced by Fiorini et al. (2024), built over $n=500$ nodes and split into $c=5$ classes. Each class contains 30 random intra-class hyperedges, while inter-class directed hyperedges, consisting of multiple tail and head nodes, are added between class pairs with sizes drawn uniformly from $\{3,\ldots,10\}$. By varying the number of inter-class hyperedges $I_{o}\in\{10,30,50\}$, we control the strength of directional connectivity. This design provides a clean benchmark to test the models’ ability to capture directionality; further details are given in Section D.4.
@@
The results in Table 3 clearly demonstrate the advantage of our models DSHN and DSHNLight over existing baselines. Classical undirected hypergraph methods are unable to capture the directional structure that dominate these benchmarks, and as a result their performance is limited. Directed methods such as GeDi-HNN and DHGNN achieve stronger results, confirming the importance of explicitly incorporating directionality into the convolutional process. Yet, DSHN and DSHNLight, which provide a principled and more expressive treatment of directional structure, yield consistent improvements across all synthetic datasets, outperforming the strongest directed baselines by up to 18 percentage points and reaching 99.04% accuracy on the third synthetic dataset—this highlights the expressive power that the notion of Directed Hypergraph Cellular Sheaves unlocks.
@@ 6 Conclusion and Future Works
We introduced the concept of Directed Hypergraph Cellular Sheaves for directed hypergraphs and derived the corresponding Directed Sheaf Hypergraph Laplacian, which we integrated into our proposed framework DSHN. By encoding hyperedge direction via a topology-aware complex-valued inductive bias, our method naturally accommodates both directed and undirected hypergraphs while also unifying and generalizing several operators from the graph and hypergraph learning literature. Across a broad set of benchmark datasets, DSHN consistently outperforms methods from both the directed and undirected hypergraph learning literature. As future work, a natural step forward is to evaluate our framework on larger and natively directed hypergraph datasets such as protein-protein interaction networks to further test the scalability and expressivity of the method, possibly employing Language Models (LMs) to generate features. Finally, an intriguing direction is to make the charge parameter $q$ directly learnable, allowing each layer to adapt its diffusion process dynamically.
Acknowledgement of Support
Antonio Purificato, Federico Siciliano and Fabrizio Silvestri acknowledge projects FAIR (PE0000013), under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU, and project NEREO (Neural Reasoning over Open Data), funded by the Italian Ministry of Education and Research (PRIN) Grant no. 2022AEFHAZ. Stefano Coniglio’s work was partially supported by the European Union under Next Generation EU — the Italian National Recovery and Resilience Plan (PNRR), PRIN 2022 PNRR (project code P20227CTY3, CUP D53D23018800001), project title ”HEXAGON: Highly-specialized EXact Algorithms for Grid Operations at the National level”.
Reproducibility Statement
We provide all the necessary information to facilitate the reproducibility of our results. Our code repository code can be found here. The README contains all that is needed to set up the Python environment and run the experiments with the different configurations. Further details on the Experimental Setup can be found in Appendix D.
Ethics Statement
All datasets employed in this work are publicly available for research and contain no personally identifiable information or harmful content (see Section D.4 for further details). The methods introduced in this paper have a societal impact comparable to that of other graph neural networks.
LLM Usage Statement
All technical content presented in this paper is entirely our own work, with LLMs serving only as an editorial tool. No scientific content or research findings were generated using an LLM.
@@ A.1 Spectral Properties
The following Lemma (which we state for clarity even if it is not reported in the paper as a lemma) derives the expression of our proposed Laplacian matrix $\mathbf{L}^{\vec{\mathcal{F}}}$ when applied as a linear operator on a signal:
@@
Let $x\in\mathbb{C}^{nd}$ be a complex-valued signal. Component-wise, the application of $\mathbf{L}^{\vec{\mathcal{F}}}$ to it and to its normalized counterpart reads:
$\bigl(\mathbf{L}^{\vec{\mathcal{F}}}(\mathbf{x})\bigr)_{u}=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{x}_{v}\right).$
$\bigl(\mathbf{L}^{\vec{\mathcal{F}}}_{N}(\mathbf{x})\bigr)_{u}=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\Bigl(\mathbf{D}_{u}^{-\frac{1}{2}}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\Bigr)\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{D}_{u}^{-\frac{1}{2}}\mathbf{x}_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{D}_{u}^{-\frac{1}{2}}\mathbf{x}_{v}\right).$
@@
We start by applying the definition of the $\mathbf{L}^{\vec{\mathcal{F}}}$ component-wise as in Eq. 3:
$\bigl(\mathbf{L}^{\vec{\mathcal{F}}}(\mathbf{x})\bigr)_{u}$ $=$ $\sum_{v\in V}(\mathbf{L}^{\vec{\mathcal{F}}})_{uv}\,\mathbf{x}_{v}$
$=$ $\sum_{e:\,u\in e}\left(1-\frac{1}{\delta_{e}}\right)\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\sum_{e:\,u\in e}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{x}_{v}$
$=$ $\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\left((\delta_{e}-1)\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{x}_{v}\right)$
$=$ $\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\left((\delta_{e}-1)\,\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{x}_{v}\right).$
Finally, notice that the coefficient $\delta_{e}-1$ is exactly the number of vertices in $e$ different from $u$. Thus, the term $(\delta_{e}-1)\,\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{x}_{u}$ can be written as a sum of $\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{x}_{u}$ over all $v\in e,v\neq u$. Substituting this back, we obtain:
$\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\left(\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{x}_{v}\right)$
$\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{x}_{v}\right).$
@@
The claim follows rather directly since, as it is not hard to see, $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ is Hermitian by construction. ∎
@@
By definition of the Dirichlet energy as the quadratic form associated with $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$, we have:
$\mathcal{E}_{N}(\mathbf{x})=\mathbf{x}^{\dagger}\mathbf{L}^{\vec{\mathcal{F}}}_{N}\mathbf{x}=\sum_{u\in V}\mathbf{x}_{u}^{\dagger}\bigl(\mathbf{L}^{\vec{\mathcal{F}}}_{N}(\mathbf{x})\bigr)_{u}.$
By substituting for $(\mathbf{L}^{\vec{\mathcal{F}}}_{N}(\mathbf{x}))_{u}$ (see the previous lemma), we have:
$\mathcal{E}_{N}(\mathbf{\mathbf{x}})$ $=\sum_{u\in V}\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr)^{\dagger}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}\right).$
@@
$\mathcal{E}_{N}(\mathbf{x})$ $=\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{u\in e}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}-\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}$
$=\sum_{e\in E}\frac{1}{\delta_{e}}\left(\sum_{u\in e}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\bigl\|\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr\|_{2}^{2}-\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}\right).$
Since $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ is Hermitian, the second inner summation can be rewritten as:
$-\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}=$
$-\sum_{\begin{subarray}{c}u,v\in e\\ u<v\end{subarray}}\left((\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}+\bigl(\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}\bigr)^{\dagger}\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\right)=$
$-\sum_{\begin{subarray}{c}u,v\in e\\ u<v\end{subarray}}2\,\Re\left[\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}\right]=$
$-\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\Re\left[\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}\right].$
@@
$\mathcal{E}_{N}(\mathbf{x})$ $=\frac{1}{2}\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\Bigl(\|\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\|_{2}^{2}+\|\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}\|_{2}^{2}-2\,\Re\left[\bigl(\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}\bigr)^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}\right]\Bigr).$
@@
$\mathcal{E}_{N}(\mathbf{x})=\frac{1}{2}\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}u,v\in e\\ u\neq v\end{subarray}}\Bigl\|\vec{\mathcal{F}}_{u\trianglelefteq e}\mathbf{D}_{u}^{-\tfrac{1}{2}}\mathbf{x}_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\mathbf{D}_{v}^{-\tfrac{1}{2}}\mathbf{x}_{v}\Bigr\|_{2}^{2}.$